OpenStudy (anonymous):

please help trying to study and prepare myself for this alg 2 test on wensday

OpenStudy (anonymous):

\[2(x-50) ^{1/3} = -10\]

OpenStudy (anonymous):

first you should divide the equation by 2 to get (x-50)^1/2=-5 then square the two side of the equation to get x-50=25 & transfer -50 to other side to get x=75